OpenStudy (anonymous):

the population P (in thousands) of a town can be modeled by P=2|t-6|+4, where t=0 represents 1990. During which 2 years does the town have a population of 8000

OpenStudy (unklerhaukus):

\[8=2|t-6|+4\] solve for \(t\)

OpenStudy (anonymous):

you can't just give me the answer?

OpenStudy (anonymous):

@bohotness

OpenStudy (bohotness):

unkle i think yuo got this

OpenStudy (unklerhaukus):

take-away 4 form both sides, then divide both sides by 2

OpenStudy (anonymous):

take away 4 from 2 and 6?

OpenStudy (unklerhaukus):

\[8\quad\ \ \ =2|t−6|+4\\8\color{brown}{-4}=2|t−6|+4\color{brown}{-4}\]

OpenStudy (unklerhaukus):

... = ...

OpenStudy (anonymous):

........

OpenStudy (unklerhaukus):

..

OpenStudy (anonymous):

please just give me the answer

OpenStudy (unklerhaukus):

do the easy steps

OpenStudy (anonymous):

nvm im just gonna put something random and delete my account this site is dumb

OpenStudy (unklerhaukus):

i'm telling you how to solve it, but you have to do the work,

OpenStudy (unklerhaukus):

8-4 = ? 4-4 = ??